Consider an economy that is characterized by the following equations:



C = 150 + 0.65(Y - T) - 200r



T = 100 + 0.2Y I = 200 - 200r



G = 500



X = 100



IM = 150 + 0.1(Y - T) - 100r



L = - 25 + 0.5Y - 500r



M = 133, 200



P ^ (SR) = 120



(Consumption)



(Taxation)



(Investment Demand)



(Government Expenditure)



(Exports)



(Imports)



(Money Demand) (Money Supply)



(Short-Run Price Level)



Answer each of the following questions. In your answers, be sure to state any assumptions that you impose and provide an explanation.



Derive the AD and SRAS curves.



Solve for the short-run equilibrium in the AD-SRAS model. Is your solution the same as in part 3 above? Why or why not?



Is the fiscal multiplier in this economy larger or smaller than if the asset market were not accounted for in the model? Briefly explain.



True or false? The aggregate demand curve is downward-sloping because the demand for goods and services increases as the price decreases. Briefly explain.

1
Expert's answer
2022-02-21T13:30:39-0500



1.) AD=C+I+G+Nx

AD=150+0.65(Y-T)-200r+200-200r+500+100-150+0.1(Y-T)-100r

AD=800+0.75(Y-T)-500r

2.) Short run equilibrium

AD=SARS

3.) If asset market was included, fiscal multiplier will be large.

4.) True. This is because of the inverse relationship between output and price.